# Beam Splitter Physics and Testing Bell Inequalities with Phase Retarders

### v2

@@ Abstract
Now, first let us verify that the effect of the AC phase cannot be observed with a source of the electrical field linear charge distribution placed in the setup given in Figure 2. With the selection of initial state (4) for the spatially-correlated (or spin-entangled) particle pair, the final state of the system is $\left|\psi_{out}\right\rangle^{\prime}=({BS_{L}}\otimes{BS_{R}})({T_{L}}\otimes{T_{R}})(P_{L}\otimes P_{R})\left|\psi_{in}\right\rangle$. Here, $T_{L(R)}$ is the quantum topological phase operator,
@@
Second, in order to observe such effects, the physical/quantum mechanical processes in Fig. 3 can be improved by adding linear charge distributions (See Fig. 5.b). Similarly, by choosing (4) as the initial state, the output state of the system is $\left|\psi_{out}\right\rangle=({BS_{L}}_{2}\otimes{BS_{R}}_{2})({T_{L}}\otimes{T_{R}})(P_{L}\otimes P_{R})({BS_{L}}_{1}\otimes{BS_{R}}_{1})\left|\psi_{in}\right\rangle$. By neglecting the overall phase factor, and taking into consideration (3) and (18), the final state of the system can be rewritten explicitly as ($\lambda=\lambda_{L}-\lambda_{R}$),
$|\psi_{out}\rangle$ $=$ $\frac{1}{\sqrt{2}}[(-\sin{\frac{\vartheta_{L}}{2}}|{D_{0}}^{\prime}\rangle+\cos{\frac{\vartheta_{L}}{2}}|{D_{1}}^{\prime}\rangle)\otimes(\cos{\frac{\vartheta_{R}}{2}}|D_{0}\rangle+\sin{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)$ (23)
$-e^{-2i\mu\lambda}(\cos{\frac{\vartheta_{L}}{2}}|D_{0}^{\prime}\rangle+\sin{\frac{\vartheta_{L}}{2}}|D_{1}^{\prime}\rangle)\otimes(-\sin{\frac{\vartheta_{R}}{2}}|D_{0}\rangle+\cos{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)]$
The phase factor in (20) appears in the expectation value expressions in contrast to (19), and it allows the correlation function S to be controlled by a physical quantity placed in the singular region without any classical interaction. To demonstrate this, we first find the joint-detection probabilities (or analogously the joint-spin measurement probabilities for arbitrary $\bm{\hat{\tilde{n}}}$ directions) which reveals the statistical behavior of quantons,
@@
$P(D_{0}^{{}^{\prime}},DD_{1})\equiv P(\alpha\Uparrow_{\tilde{n}},\beta\Downarrow_{\tilde{n}})$
$=\frac{1}{4}[1+\cos{\vartheta_{L}}\cos{\vartheta_{R}}+\sin{\vartheta_{L}}\sin{\vartheta_{R}}\cos{(2\mu\lambda)}]$ (24)
Then, by introducing observables $\mathcal{O}^{L(R)}$ as in equation (11), the expectation values are revealed,
@@
Last, by substituting (22) into (13), with the certain chosen angles the correlation function S is obtained as,
@@
Equation (20) illustrates how the S function depends on the physical quantities in the singular region. Accordingly, for specific angle selections, the S function can be controlled by the topological phases. In the case of a linear charge distribution $\lambda=0$, it reduces to $S=2\sqrt{2}$. Moreover, these are identical to the results of [7] obtained in a spin-correlated AC system and to the results of [Bertlman], where the Berry phase is studied in the case of confined azimuthal angles. The fact that these results are analogous to spin measurements based on detection probabilities shows the importance of BS-P-BS systems and BS applications in emerging quantum technologies.
Conclusion.—The main aim of this study is to analyze the characteristics arising from the use of BS in combination with P. In BS-P-BS systems, we first demonstrate that the detection probabilities of quantons depend on the phase of P, and these probabilities are equivalent to the probabilities of spin measurement in arbitrary directions in analog setups. Here, the phase of P is equivalent to the angle of the direction in which the spin is measured. Then, we propose two possible EPR-Bohm-type experimental setups for developing this use of BS for two-quanton systems.
In the first scenario, spatially-correlated quantons generated from the same source pass directly through P and reach the BSs. The (7) probabilities are related to the joint-spin measurement probabilities in EPR-Bohm type setups (with a $\frac{\pi}{2}$ phase difference as in the diffraction experiments). On the other hand, the effects of the AB phase on the spatially-correlated quanton system were studied by Silverman in a similar proposed setup without P, where the phases are predicted to affect the probabilities of detection (see [10] equations 4-5). However, the paths of the moving quantons do not form a closed trajectory, and it is not a suitable setup for observing the topological phase effects. Moreover, since the AB (and Dual AB) phases are independent of the spin orientations of the incoming quantons, the effects of these phases on spin-entangled states do not arise in the expectation value expressions [26-28]. In conclusion, although the results of Silverman’s work are inspiring, the relevant probabilities cannot be obtained using this setup, and it needs to be modified by the P and spin-dependent AC (or HMW) phase [23-25].
Secondly, a BS-P-BS system is proposed in which two spatially-correlated quantons produced from the same source follow closed trajectories (as in interference experiments). The results demonstrate that the joint-detection probabilities are equal to the probabilities of joint-spin measurements for spin-entangled quantons in EPR-Bohm-type setups. Accordingly, the spin-up and spin-down probabilities for measurements in arbitrary directions in spin-entangled systems are equal to the probabilities of detection by the up-placed detector and the down-placed detector for the determined phase of retarders analogously. These are crucial results and allow us to test the BI. Thus, one can reach the expectation value expressions and the correlation function S, which depends on detection probabilities and the phase of the retarder by introducing an observable. Here, the violation of BI can be followed by controlling P.
Moreover, in contrast to the previous setup, the correlated BS-P-BS system is appropriate for investigating the effects of topological phases. In this regard, the BS-P-BS configuration is reconsidered in the presence of the spin-dependent AC (or HMW) phase. Now, the obtained probabilities depend on the physical quantity in the singular region as well as the phase of the retarder and identical to reference [7], where the effects of topological phases are examined in the framework of entirely relativistic quantum mechanics using spin-entagled systems in 2+1 dimensions. Furthermore, geometric phases that occur in both classical and quantum systems adiabatically evolving through a closed cycle defined in the configuration space of its external parameters are special cases of topological phases [29-31]. Hence, it is necessary to verify the obtained results for the study by Bertlman, where the effects of geometric phases on the entanglement quantum states are investigated [32-34]. Indeed, it is understood that the probability expressions for [32] results will be reduced to expression (21) by confining the azimuthal angles for 2+1 dimensions. This is also a correction to [10] and gives the probability expressions intended to be found. From this point, the intended results can be achieved by choosing the matrix elements,
@@
for r and t coefficients analogous to the phase of the retarder. Although (24) implies that the results can be obtained using SBSs that are not symmetric, this would be unusable in practice. The necessity of choosing a different SBS each time does not satisfy the spin measurement analogy. Last, the proposed setup is utilized as a new testing method for BI. The correlation function becomes dependent only on the physical quantity in the singular region with the phases (angles) that are maximally violated by quantum mechanical expectation values. Consequently, the non-local properties of quantum mechanics can be tested simply using linear charge distribution and magnetic dipole moment carrier chargeless quantons.
The author would like to thank Prof. Ali Ulvi Yilmazer, Prof. Anatoli Polkovnikov, Dr. Melik Emirhan Tunalıoğlu and Dr. Tatsuhiko Ikeda for their valuable contributions. This work is supported by TUBITAK-BIDEB 2219 Research Fellowship.

### v3

@@ 3 Testing Bell-CHSH inequalities using topological phases
Now, first let us verify that the effect of the AC phase cannot be observed with a source of the electrical field linear charge distribution placed in the setup given in Fig. 2. With the selection of initial state  (4) for the spatially-correlated (or spin-entangled) particle pair, the final state of the system is $\left|\psi_{out}\right\rangle^{\prime}=({BS_{L}}\otimes{BS_{R}})({T_{L}}\otimes{T_{R}})(P_{L}\otimes P_{R})\left|\psi_{in}\right\rangle$. Here, $T_{L(R)}$ is the quantum topological phase operator,
@@
Figure 4: (a) Schematic representation of the AC phase in 2+1 dimensions resulting from the motion of magnetic dipole moment carrier chargeless particles (or HMW phase in the case of $\bm{\tilde{E}}\longrightarrow\bm{\tilde{B}}$ and $\mu\longrightarrow d$) in a closed trajectory around a linear charge distribution. (b) A proposal for a gedanken experimental setup to investigate the effects of topological phases on entangled quantum states and to observe the violations of BI using physical quantities in the singular region to test the non-local properties of quantum mechanics.
@@
Second, to observe such effects, the physical/quantum mechanical processes in Fig. 3 can be improved by adding linear charge distributions (See Fig. 4.b). Similarly, by choosing  (4) as the initial state, the output state of the system is $\left|\psi_{out}\right\rangle=({BS_{L}}_{2}\otimes{BS_{R}}_{2})({T_{L}}\otimes{T_{R}})(P_{L}\otimes P_{R})({BS_{L}}_{1}\otimes{BS_{R}}_{1})\left|\psi_{in}\right\rangle$. By neglecting the overall phase factor, and taking into consideration  (2),  (3) and  (17), the final state of the system can be rewritten explicitly as ($\lambda=\lambda_{L}-\lambda_{R}$),
$|\psi_{out}\rangle$ $=\frac{1}{\sqrt{2}}[(-\sin{\frac{\vartheta_{L}}{2}}|{D_{0}}^{\prime}\rangle+\cos{\frac{\vartheta_{L}}{2}}|{D_{1}}^{\prime}\rangle)\otimes(\cos{\frac{\vartheta_{R}}{2}}|D_{0}\rangle$
$+\sin{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)-e^{-2i\mu\lambda}(\cos{\frac{\vartheta_{L}}{2}}|D_{0}^{\prime}\rangle+\sin{\frac{\vartheta_{L}}{2}}|D_{1}^{\prime}\rangle)$ (19)
$\otimes(-\sin{\frac{\vartheta_{R}}{2}}|D_{0}\rangle+\cos{\frac{\vartheta_{R}}{2}}|D_{1}\rangle)]$
The phase $(2\mu\lambda)$ in  (3) appears in the expectation value expressions in contrast to  (18), and it allows the correlation function S to be controlled by a physical quantity placed in the singular region without any classical interaction. To demonstrate this, we first find the joint-detection probabilities (or analogously the joint-spin measurement probabilities for arbitrary $\bm{\hat{\tilde{n}}}$ directions), which reveals the statistical behavior of quantons,
@@
$P(D_{0}^{{}^{\prime}},D_{1})\equiv P(\alpha\Uparrow_{\tilde{n}},\beta\Downarrow_{\tilde{n}})$ (20)
$=\frac{1}{4}[1+\cos{\vartheta_{L}}\cos{\vartheta_{R}}+\sin{\vartheta_{L}}\sin{\vartheta_{R}}\cos{(2\mu\lambda)}]$
Then, by introducing observables $\mathcal{O}^{L(R)}$ as in  (11), the expectation values are revealed,
@@
Last, upon substituting  (21) into  (2), we derive the correlation function S.
$\begin{split}S(\vartheta_{L},\vartheta_{R},\vartheta_{L}^{\prime},\vartheta_{R}^{\prime},\lambda)=&\left|-\cos\vartheta_{L}\cos\vartheta_{L}^{\prime}-\sin\vartheta_{L}\sin\vartheta_{L}^{\prime}\cos(2\mu\lambda)\right.\\ &+\left.\cos\vartheta_{L}\cos\vartheta_{R}+\sin\vartheta_{L}\sin\vartheta_{R}\cos(2\mu\lambda)\right|\\ &+\left|-\cos\vartheta_{L}^{\prime}\cos\vartheta_{R}^{\prime}-\sin\vartheta_{L}^{\prime}\sin\vartheta_{R}^{\prime}\cos(2\mu\lambda)\right.\\ &-\left.\cos\vartheta_{R}\cos\vartheta_{R}^{\prime}-\sin\vartheta_{R}\sin\vartheta_{R}^{\prime}\cos(2\mu\lambda)\right|\end{split}$ (22)
Thus, the correlation function (analogous to Bell angles) for which the retarder phases are maximally violated by the quantum mechanical expectation values is obtained as follows:
@@
Figure 5: Maximal values of the S function can be explicitly controlled by the AC-HMW phases. Accordingly, without loss of generality, one of the retarder phases can always be eliminated $\vartheta_{L}=0$. By calculating the extremum conditions of the function (23) for the other phases of retarders, one can consider the maximal $S$ values for $\mu\lambda$s with the solutions $\vartheta_{L}^{\prime}=\pi/2$, $\vartheta_{R}=\pm\arctan[\cos{2\mu\lambda}]$, and $\vartheta_{R}^{\prime}=\pi-\vartheta_{R}$.
This result illustrates how the correlation function depends on the physical quantities in the singular region. Accordingly, for specific angle selections, $S$ can be controlled by linear charge distributions (See Fig. 5). In the case of $\lambda=0$, it reduces to $S=2\sqrt{2}$. Besides, when the polarized dipoles are used, there exists entanglement between the spin and path of chargeless particles as its different degrees of freedom. In this case, it is physical noncontextuality rather than the locality that is tested experimentally.
The proposed setup is important for quantum communication and information applications, as the generalizations of BI serve as criteria for entanglement and separability. The nonlocal structure of AC-HMW phases facilitate operations at requisite distances and connections for BI testing, devoid of adiabatic conditions or classical interactions. The robust structure of topological phases offers the possibility of obtaining more precise results due to their resistance to environmental noise. Furthermore, in the setup utilizing spatially correlated particles (with momentum conservation),  (23) is valid for all types of quantons (bosons, fermions) that provides flexibility and suitability under diverse laboratory conditions. The use of moving particles as electrically and magnetically neutral dipole carriers is adequate for experimentation  [23; 26]. In this context, the setups with neutron interferometers, where the effects of geometric phases on entangled quantum states are examined, stand out as the most inspiring studies  [29; 49; 50; 51]. Lastly, it should be further emphasized that the results of  (23) emerge solely based on the detection probabilities of particles, analogously to spin measurements, and that Bell angles for BI testing can be controlled through phase retarders.
4 Discussion and conclusions
Bell-CHSH inequalities (BI) are crucial for testing the fundamental principles of quantum mechanics, particularly in assessing quantum entanglement, one of its most profound aspects within the framework of local realism. Generalizations of BI serve as criteria for entanglement and separability, with experimental implementations conducted in large quantum systems such as photons and atoms. On the other hand, topological AC and HMW phases emerge as excellent tools in developing quantum technologies with their nonlocal features. In this study, our primary motivation is to propose gedanken experimental setups for testing BI utilizing all types of quantons, especially spin 1/2 particles, extending beyond photons. We delve into the effects of spin-dependent AC-HMW phases on entangled quantum states, leveraging their nonlocal structure to gain an experimental advantage. Therefore, we place electromagnetic field sources (linear charge distributions) in the center of BS-P-BS systems, where the motion of particles is not allowed. Through this setup, we observe the topological phase contributions to the wave functions of the moving dipole carrier chargeless particles. Consequently, the correlation function S can be explicitly controlled by the topological phase, thereby influencing the entangled quantum states.
Now, let us have some discussion on our scheme and its possible implementations in physical systems. First, the AC phase is recognized as a special case of the Berry phase that emerges under adiabatic conditions  [52; 53]. In this regard, the research conducted by Bertlmann is significant  [29]. In this work, the authors explore the effects of the Berry phase generated by implementing an adiabatically rotating magnetic field into one of the paths of magnetic dipole carrier entangled neutrons moving in opposite directions. They use spin-echo method to eliminate phases that arise due to the system dynamics during the motion of the particles and perform with neutron interferometers. Since their results (See equations (20-22) in  [29]) should be reduced to our  (23) with Bell angles by restricting the azimuthal angles, it can be reconsidered for the proposed setup in Fig. 4. In contrast to  [29], our proposed setup offers a significant advantage by utilizing AC-HMW phases, which possess non-local features, without requiring an adiabaticity condition. In this regard, it also clearly reveals the relationship between Berry phase and its special cases AC-HMW phases. Moreover, at this stage, a discussion on noncontextuality needs to be addressed. Noncontextuality requires that observables yield a value that does not depend on the experimental context; results of measurements must be independent of other simultaneous measurements. When the AB effect is considered in two dimensions, the solenoid can be regarded as a dipole and is inherently polarized in the z-direction. Hence, the identity relation between the AB and AC phases holds when AC dipoles are polarized (for $s=\pm 1$). In the proposed setup, entanglement occurs between the spin and paths of neutral particles. This suggests experimentally testing physical noncontextuality rather than locality [48-51]. Thus, we use the BI to test noncontextual hidden variable theories.
Second,  (23) is also consistent with the findings of [19]. In this study, the authors delve into the AB and AC effects within the framework of fully relativistic quantum mechanics in 2+1 dimensions, demonstrating the ability to control the correlation the correlation function for BI through the AC phase. To achieve this, they leverage spin measurement probabilities and associated expectation value expressions. Here, we utilize the probabilities of particle detection analogously to spin measurements and use phases of the retarders analogous to Bell angles in the proposed setup. This illustrates the application of Mach-Zender interferometers modified with phase retarders, where the probabilities of detection by the detectors serve as analogs to spin measurements. As a result, this offers an example of conducting entanglement experiments using all kinds of quantons, not limited to photons. Thus, we pave the way for experimental exploration of the effects of phases on entangled quantum states, verification of the spin independence of the AB phase, and examination of the duality and identity relationships between phases, considering the HMW phase as well. Such endeavors are important for quantum communication and information applications.
Third, as a clear example of AC-HMW duality, it reveals the nature of AB-type quantum mechanical effects. It is conceivable that in such an configuration the best known and earliest example of topological phases, the AB phase (which is also the case for the fully dual DAB phase), could be used at the first instance. However, the interaction Hamiltonian describing the dynamics of the AB system and the transformed wavefunction are given by,
where e is the electron charge, $\bm{A}$ is the vector potential, and $\Phi$ is the magnetic flux. The phase is independent of the spin orientation of the particles (s), and disappears in the expectation value expressions. The results are reduced to the equations  (2) and  (2) even in the presence of the solenoid in the case of using AB phase in the proposed setups contrarily done in Silverman’s study  [35]. Moreover, although the setup in Fig. 2 works for all types of quantons (bosons, fermions, etc), it is not suitable to investigate the effects of topological phases on entangled quantum states even if we consider the spin-dependent AC-HMW phases. Hence, the probabilities in  (7) are related to the joint-spin measurement probabilities in EPR-Bohm-type setups with phase difference $\frac{\pi}{2}$, and since no real closed orbit is formed (single-slit analogy). The configuration in Fig. 3 is well suited to study the effects of nonlocal topological phases (double-slit analogy) and it leads to construction of Fig. 4  [34].
Last, although the study can be performed using polarizing BSs, using a different BS each time to perform the spin measurement analogy is not practical for experimental applications. Nevertheless, for the completeness of the study, it can be realized by choosing coefficients r and t for  (1) as,
@@ 4 Discussion and conclusions
The proposed setup, constructed with spatially correlated particles (via conservation of momentum) analogously to singlet states, does not allow for the examination of mixed states. Nevertheless, conducting such a study is imperative for achieving a thorough understanding of the influence of geometric and topological phases on quantum mechanical states. In this regard, we intend to pursue this line of research in future studies. On the other hand, we believe that modifying the proposed setup to incorporate four entangled quantons will enable the testing of Leggett’s inequality, with the aim of exploring the incompatibility between nonlocal realism and quantum mechanics using AC-HMW phases  [30]. Similarly, it could serve as an experimental setup suggestion for the study with four entangled quantons  [28].
Acknowledgements
The author would like to thank Prof. Ali Ulvi Yilmazer, Prof. Abdullah Verçin, Prof. Anatoli Polkovnikov, Dr. Melik Emirhan Tunalıoğlu, Dr. Tatsuhiko Ikeda and Mary Lynch for their valuable contributions. This work is supported by TUBITAK-BIDEB 2219 Project.